PLEASE HELP ME W/ CHANGING COMPLEXPERCENTS TO DECIMALS LIKE {1 21/25} PLEASE SHOW ME HOW THE WORK IS DONE
Answer by josmiceli(19441)

per cent means per 100
cent is the old Latin word for 100
so when you see "%" think "/100"
57% = 57 / 100 = .57
each time youi move the decimal point to the left, you divide by 10
move it twice, and yoou divide by 10x10 or 100
-----------
how do you convert a decimal to a fraction?
just keep moving the decimal point to the right until you run out of places
dividing by 10 each time
.54821357226 =
5.4821357226 /10 =
54.821357226 / 100 =
548.21357226 / 1000 =
5482.1357226 / 10000 =
54821.357226 / 100000 =
548213.57226 / 1000000 =
5482135.7226 / 10000000 =
54821357.226 / 100000000 =
548213572.26 / 1000000000 =
5482135722.6 / 10000000000 =
54821357226 / 100000000000
--------
so .54821357226 = 54821357226 / 100000000000
and you converted a decimal to a fraction
a simple one like .95 =
9.5 / 10 =
95 / 100
divide top and bottom by 5
95/100 = 19/20
(the LCD is 5)